Subject: obs log 20000305 (mjd 51609) Date: Mon, 6 Mar 2000 06:35:31 -0700 (MST) From: hayes@apo.nmsu.edu (Jeffrey Hayes) [mail to: obs-report-25m@apo.nmsu.edu include MJD in subject] Sunday Mar 05 2000 Apache Point Observatory SDSS 2.5m Observing Log (MJD 51609) OBSERVING TEAM: =============== Jeffrey Hayes (APO) Shift 2 Ellyne Kinney (APO) Shift 2 Scot Kleinman (APO) Shift A Atsuko Nitta Kleinman (APO) Shift A Bing Chen (JHU) PT VISITORS: ========= Andrew Connolly (UP) Two Pies (Cldcrft) THE PLAN: ========= Spectro all night unless conditions change and look favorable for the remainder of the night. SYNOPSIS FOR NIGHT: =================== One FASTT plates, lots of clouds. TELESCOPE STATUS: ================= DIMM: ----- When open, from 1.0 to 2.1 arc-sec. INTERLOCKS: =========== None bypassed. SOFTWARE USED: ============== IOP/SOP: v3_10_1 Watcher: v2_6_7 MCP: v4_0_1 WEATHER SYNOPSIS (see below for hourly data): ============================================= Never photometric; rarely spectroscopic. GENERAL COMMENTS (see Problems below for details): ================================================== First off: cart 3 plate 316. During pre-twilight, we tried to take a spectro bias with the default poolInit. Same problems as last night. We tried to recover sd1 with various incarnations of endStare and iccExec $active readoutAbort. This didn't work, so we rebooted the crate and re-issued our own poolInit (1 1 9500000 9500000) and the next bias worked fine. While waiting for twilight, we exercised the new (3_10_1) version of sop and verified preCalib, postCalib, and diamondPoint were working. We loaded a full set of tapes in order to run some imager biases during the night. ID1-1 came up with the write protect light (again) even though the tape wasn't write-protected. We inserted a new tape and it loaded fine. We had ID2 problems throughout the night. By 21:20, we were offToField on cart 4 plate 266. We closed due to increasing clouds and a dome that's not easy to close and a telescope that's not easy to protect in an emergency around 22:42. At 01:00 we opened again and swapped cartridge 4/266 for cartridge 8/292 By 01:30 we were back on the sky. We did preCalibs and found the focus to be 1450. We also started guiding. We started science at 01:50. 10 minutes into the first integration, the focus was the same as our initial one: 1450. During the 4th science integration the focus was changed to 1400. We declared this plate done at 03:40. At 03:30, we switched to cartridge 1/plate 304. Precalibs done by 04:00. We did not see any stars after the initial slew to the field, but we did successfully do an fk5InFiber, and after making the offsets sticky, slewed to the field again where we found guide stars. Focus set to 1450. First science integration started at 04:20. 3 subsequent science integrations, the diamond point and the postCalib took us to twilight. At this point we declared the plate done and closed up. Ellyne ran "endNight 51609 -doSpectra" which worked. (And there was great rejoicing....) DATA ACQUISITION (see IOP/SOP output below): =========================================== IMAGING: -------- Run 1243: bias drift started at 19:30 -ended at 19:55 or so due to archiver problems. (See below.) SPECTROSCOPY: ------------- 2787 2788 2789 60s Science on plate 316. 2792 - 2795 diamondPoints for plate 316. 2797-2802 -- biases 2803-2805 -- darks 2808-2810 -- 900s each Science on plate 292 2813 -- 900s Science on plate 292 2814-2815 -- diamondPoints for plate 292 2826-2829 -- 900s each Science on plate 304 2830-2831 -- diamondPoints for plate 304 MIRRORS: ======== PRIMARY: -------- MIGS SHOULD BE Axial A= 1.9315 1.9315 Axial B= 0.5565 0.5570 Axial C= 0.5559 0.5570 Transverse D= 0.4084 0.4075 Lateral E= 1.8985 1.8980 Lateral F= 0.3989 0.3990 GALILS Commanded: 000004850, -000002350, 000000100, -000009500, 000000050, -00000100 Actual: 000004850, -000002350, 000000100, -000009500, 000000050, -000000100 SETMIR VALUES primDesOrient: primOrient: SECONDARY: ---------- Focus: 1300 Temp: 1.2C (t_collimate not yet active) Altitude: 35 MIGS Axial A= -0.0609 Axial B= -0.0459 Axial C= -0.0649 Transverse D= 0.1040 GALILS Commanded: 001517500, 001504650, 001494350, -000006950, -000011550 Actual: 001534660, 001522654, 001510578, -000006950, -000011550 SETMIR VALUES secDesOrient: secOrient: LN2 SCHEDULE: ============= Filled spectro dewars around 18:30. And again at 06:00. PROBLEMS: ========= IOP tapeLoad ------------ Sun Mar 05 19:18:42 MST 2000: tapeWaitFor reports success Waiting for command response from id2 (crate 1, ICC 1), archiver server (client 2) Waiting for command response from id2 (crate 1, ICC 1), archiver server (client 2) Sun Mar 05 19:19:19 MST 2000: IOP Not ready (blew up on pad): tapeMount: Request sense failure Failed to receive exabyte command completion status invoked from within "iccExec $tapeList archReadSense 2" invoked from within "set sense2 [iccExec $tapeList archReadSense 2]..." (procedure "tapeRewind" line 12) invoked from within "tapeRewind $tapeList " invoked from within "if {$doRewind} { tapeRewind $tapeList }" (procedure "tapeLoad" line 5) invoked from within "tapeLoad $tapeList $doRewind $doUnload" im> Modu="mir_Adjust"; Text="Cannot compute PrimOrient" A repeated initNight -tapeLoad worked fine. IOP ID2 ------- Mar 05 19:29:19 sdssid2 dArch dsc_I_frameArchived Frame 1243 4.12 was archived to tapes JL2041 (21) and JL2116 (21) Mar 05 19:29:19 sdssid2 dArch dsc_I_frameArchived Frame 1243 6.22 was archived to tapes JL2041 (22) and JL2116 (22) Mar 05 19:29:22 sdssid2 dArch dsc_E_trcErr dscArch_CheckPrevOp: rstatus = 0x8001421c, drive 1 Mar 05 19:29:22 sdssid2 dArch dsc_E_archShutDown1 Archiver shutting down first drive Mar 05 19:29:22 sdssid2 dArch dsc_E_archWrFail Tape write error while archiving frame 1243 8.32 Mar 05 19:29:24 sdssid2 dArch dsc_E_trcErr dscArch_CheckPrevOp: rstatus = 0x8001421c, drive 2 Mar 05 19:29:24 sdssid2 dArch dsc_E_archShutDown2 Archiver shutting down second drive Mar 05 19:29:24 sdssid2 dArch dsc_E_archWrFail Tape write error while archiving frame 1243 8.32 Mar 05 19:29:24 sdssid2 dArch dsc_E_archStStopped Stopped - both drives are disabled Mar 05 19:29:24 sdssid2 dArch TEXTONLY Both tape drives disabled, so stopping archiver Mar 05 19:29:25 sdssid2 dArch dsc_E_archStStopped Stopped - both drives are disabled Mar 05 19:29:25 sdssid2 dArch TEXTONLY Both tape drives disabled, so stopping archiver Mar 05 19:29:25 sdssid2 dArch dsc_E_archStStopped Stopped - both drives are disabled Mar 05 19:29:25 sdssid2 dArch TEXTONLY Both tape drives disabled, so stopping archiver ... Mar 05 19:29:33 sdssid2 dArch dsc_E_archStStopped Stopped - both drives are disabled Mar 05 19:29:33 sdssid2 dArch TEXTONLY Both tape drives disabled, so stopping archiver Mar 05 19:29:33 sdssid2 dArch TEXTONLY Archiver: both drives bad in auto-recover, disabling archiver Mar 05 19:29:33 sdssid2 dArch TEXTONLY Archiver: auto-recover rewinding bad drive 1 Mar 05 19:29:33 sdssid2 dArch dsc_E_archStStopped Stopped - both drives are disabled Mar 05 19:29:33 sdssid2 dArch TEXTONLY Both tape drives disabled, so stopping archiver ... Mar 05 19:31:03 sdsshost IDA 289382 TEXTONLY NO ENABLED DRIVES on id2. Do something!!! Mar 05 19:31:03 sdsshost IDA 289382 TEXTONLY id2:1 eom=-10 remTape=30399889 bytes=17011 field=4 paused=0 rampup=1 Mar 05 19:31:03 sdsshost IDA 289382 TEXTONLY id2:2 eom=-10 remTape=30399889 bytes=17011 field=4 paused=0 rampup=1 Standard recovery procedures worked, then an accidental tapeLoad was issued and things failed again, only to be recovered again for a few frames before it failed again... WEATHER DATA: ============= Wind Time Temp F Dewp F MPH Direction Dust Sky Precip. 16:01 37 9 15 255 (WSW) 489 17:01 38 8 18 205 (SSW) 529 18:01 35 12 20 208 (SSW) 772 19:01 34 12 17 229 (SW) 674 20:01 34 13 18 216 (SW) 693 21:01 35 15 17 208 (SSW) 733 22:01 34 16 15 202 (SSW) 644 23:01 35 16 23 221 (SW) 535 clouds 00:01 35 17 21 202 (SSW) 567 clouds 01:01 34 18 20 213 (SSW) 510 clouds 02:01 34 18 22 207 (SSW) 541 03:01 35 19 16 203 (SSW) 559 DATA LOGS FROM IOP/SOP: =======================